# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes (Experiment, Analysis & Benchmark Paper)

### v3

@@ 4.1. Retrieving the candidates
The main drawback of Exact Matching is the computation has a cost that depends directly on the size of the data lake and the tables therein, and this operation must be repeated for every new query column. As we show in Fig. 10, the cost for multiple query columns quickly adds up. Another drawback is that containment is less reliable if the data lake features a lot of redundancy (e.g., YADL 50k in Fig. 2), since unrelated tables may still feature high containment.
@@
The main drawback of MinHash is its lack of a candidate ranking, which means that selecting a subset of the candidates does not guarantee that the candidates will have high containment (Figure 8a). Moreover, since MinHash returns an approximate result, it may happen that candidates that have a measured containment lower than the threshold are proposed: indeed, when the MinHash query results are compared with the results obtained from Exact Matching, MinHash produces a large fraction of False Positives. Finally, while the index can be updated when new data is added to the data lake: performance may deteriorate after a certain level of updates, especially depending on the skew of the new data.
@@
Just as it takes some of the strong suits of both MinHash and Exact, Hybrid MinHash also retains some of their disadvantages: 1) The MinHash index will degrade each time it is updated. 2) As the first filtering relies on the MinHash results, False Positives add needless cost and any candidate missed by MinHash is lost. 3) Re-ranking candidates by calculating the exact containment increases the query time substantially with respect to MinHash.
@@
We compare the simple methods described so far against Starmie (Fan et al., 2023b), a SOTA system for table unionability and join discovery. Starmie employs a contrastive learning method to train column encoders from pre-trained language models to capture contextual information within tables. Starmie performs candidate retrieval by building embeddings for each column in the data lake and for the query column; it then ranks candidates with a similarity score combining embedding cosine similarity and Jaccard Containment.
4.2. Merging base table and join candidates
In this step, the candidates that were retrieved in the previous step are combined into a new integrated table that joins the information in the base table $T$ with additional information from augmentation tables. As we highlight in Fig. 1, the merge step combines two operations that cannot be executed independently of each other: join selection, and aggregation.
Aggregation
Figure 3. Example of how a left join would duplicate rows from the base table.
Following the example in Fig. 3, a join on “Title” leads to the duplication of the two first rows in the base table, as both title values appear in two rows of the candidate table. How should we aggregate the rows to obtain the best prediction over the column “Revenue”?
We test three different aggregation strategies:
  1. (1)
    First drops all rows where the values belonging to the base table are duplicated, with the exception of the first copy.
  2. (2)
    Mean replaces all duplicated numerical (categorical) values by the mean (mode) of all values for that attribute.
  3. (3)
    DFS (Deep Feature Synthesis (Kanter and Veeramachaneni, 2015)) greedily adds new features by performing different aggregations (mean, median, count, standard deviation, etc.) over each column in the table.
Join selection
We implement and test five different join selectors. All selectors follow the fit-predict paradigm proposed by scikit-learn (Pedregosa et al., 2011), which makes extending the pipeline with new selectors very simple. All experimental runs are executed over 10 cross-validation splits: the test fraction is 0.2 in this outer cross-validation split. Each selector receives as input a pool of “$K$” candidates from a given retrieval method on a given data lake, the train and test splits of the base table, and the aggregation method to use. The train split is further split into a training (0.8) and validation set (0.2). The model is trained on the final integrated table (i.e., the base table after joining it with the candidates) produced by the selector of this training split, then the result is scored with the test split.
@@ 4.2. Merging base table and join candidates
Highest Containment Join measures the exact containment of all of the candidates, then re-ranks them by that. The candidate with the highest containment is selected and joined; ties are broken by taking one candidate at random.
@@
In all selectors that involve a join, aggregation is carried out prior to executing the join itself.  [^2] In our pipeline, aggregation is carried out before executing any join by grouping the “right table” by the join key, then applying one of the aggregation functions described above. This is to avoid materializing large joins, which have a huge cost in memory and time. Due to how aggregation is carried out, the final result is the same before and after the join, so executing it before materializing the joined table is more efficient.
4.3. Predicting with a ML model
The learning step is performed on the final integrated table on the training split returned by the join selector and evaluated on the integrated table on the test split. We evaluate two different methods: a simple linear regressor/classifier (Pedregosa et al., 2011) and the more complex method CatBoost (Prokhorenkova et al., 2018). We choose CatBoost as our “complex and effective” method because gradient boosting methods are effective at working with tables (Grinsztajn et al., 2022), and CatBoost in particular is good with categorical data, which is common in our scenario.
5. Experimental study
For our experimental campaign, we test the different sections of the pipeline over three main dimensions: prediction performance, execution time, and RAM usage. Focusing on multiple dimensions allows us to have a view of the different trade-offs across methods.
Table 3. Aggregate statistics for the tables in the data lakes, along with the size on disk of the data structures used by each retrieval methods. We report the mean (Avg) for each statistic. “N. cols” and “C. cols” refer to numerical and categorical columns respectively.
YADL YADL OpenData
Binary Base 10k 50k US
N. tables 70 30072 10059 47223 5591
Size (MB) 628.9 10051.3 5557.7 27718.1 4062.4
Tot. rows 2.01e+07 6.72e+08 2.42e+08 1.20e+09 9.57e+07
Tot. cols 1.40e+02 9.52e+04 1.27e+05 6.24e+05 1.33e+05
Avg rows 287134.3 22343.9 24014.8 25430.2 17124.5
Avg cols 2.00 3.17 12.63 13.21 23.86
Avg N. cols 0.30 0.39 3.54 3.59 11.10
Avg C. cols 1.70 2.78 9.08 9.61 12.76
Avg nulls 0.00 0.31 0.62 0.64 0.09
Data lakes
We use four YADL variants (Binary, Base, 10k, 50k) (as detailed in Section 3) and Open Data US, a data lake employed in the literature (Galhotra et al., 2023; Zhu et al., 2019, 2016); statistics are reported in Table 3.
Table 4. Statistics for the base tables.

### v6

@@ B.1 Retrieval
The main drawback of Exact Matching is the computation, whose cost depends directly on the size of the data lake and the tables therein; furthermore, the operation must be repeated for every new query column. As we show in Figure 5, the cost of querying multiple columns quickly adds up. Another drawback is that containment is less reliable if the data lake features a lot of redundancy (e.g., YADL 50k in Figure 8), since unrelated tables may still feature high containment: this problem may occur if the data lake includes multiple variants of the same table, a setting commonly occurring in industry (Halevy et al., 2016). An additional drawback of Jaccard containment is that it does not consider cardinality, so that binary tables would be considered as perfect matches while being useless in practice.
@@
The main drawback of MinHash is that it does not feature an inherent ordering of candidates. Moreover, since MinHash returns an approximate result, it may happen that candidates with an actual containment lower than the threshold are returned as False Positives. Under tight constraints, no ordering and False Positives reduce the likelihood of retrieving good candidates (Figure 16). Finally, while the index can be updated when new data is added to the data lake, performance may deteriorate after a certain level of updates, especially depending on the skew of the new data.
@@
Just as it takes some of the strong suits of both MinHash and Exact, Hybrid MinHash also shares some of their disadvantages: 1) The MinHash index will degrade each time it is updated. 2) As the first filtering relies on the MinHash results, False Positives increase cost and any candidate missed by MinHash is lost. 3) Re-ranking candidates by calculating the exact containment increases the query time substantially with respect to MinHash (Table 10).
@@
We compare the simple methods described so far against Starmie (Fan et al., 2023b), a SOTA system for table unionability and join discovery. Starmie employs a contrastive learning method to train column encoders from pre-trained language models to capture contextual information within tables. Starmie performs candidate retrieval by building embeddings for each column in the data lake and for the query column; it then ranks candidates according to the equation:
$S_{Starmie}~{}=~{}|Q\cap C_{k}^{i}|\cdot S_{C}(vec(Q),vec(C_{k}^{i}))$
where $S_{C}$ is the cosine similarity between the embeddings of the query column $Q$ and candidate $C_{k}^{i}$.
B.2 Selection
Each selector receives as input a pool of “$K$” candidates from a given retrieval method on a given data lake, the train and test splits of the base table, and the aggregation method to use. The train split is further split into a training (0.8) and validation set (0.2). The model is trained on the final integrated table (i.e., the base table joined with the candidates) produced by the selector of this training split, then the result is scored with the test split.
@@ B.2 Selection
Highest Containment Join ranks candidates by exact containment, after measuring it for each candidate. The top-$1$ candidate is joined; ties are broken by taking one candidate at random.
@@
In all selectors that involve a join, aggregation is carried out prior to executing the join itself. In our pipeline, aggregation is carried out before executing any join by grouping the “right table” by the join key, then applying one of the aggregation functions described above. This is to avoid materializing large joins, which have a huge cost in memory and time. Due to how aggregation is carried out, the final result is the same before and after the join , so executing it before materializing the joined table is more efficient.
All join selectors follow the fit-predict paradigm proposed by scikit-learn (Pedregosa et al., 2011), which simplifies extending the pipeline with new selectors.
B.3 Aggregation
Figure 7: Example of how a left join would duplicate rows from the base table.
Following the example in Figure 7, a join on “Title” leads to the duplication of the two first rows in the base table, as both title values appear in two rows of the candidate table. How should we aggregate the rows to obtain the best prediction over the column “Revenue”? We test three different aggregation strategies:
  1. 1.
    Any selects one row from each group, without considering the order.
  2. 2.
    Mean replaces all numerical (categorical) values by the mean (most frequent) of all values for that attribute.
  3. 3.
    DFS (Deep Feature Synthesis (Kanter & Veeramachaneni, 2015)) greedily adds new features by performing different aggregations (mean, median, count, standard deviation, etc.) over each column in the table.
Appendix C Experimental results
Figure 8: Exact containment measured over the top 100 join candidates for each base table (represented by a colored line) on each data lake. The x-axis reports the rank, rank 30 is the retrieval cutoff we use in our experiments.
C.1 Choosing which baselines to use
Including recent publications often turned out to be very difficult, as the code was either unavailable, or not suited to extensive experiments on a data lake of the size that we investigate in our work. We worked on including Starmie (Fan et al., 2023b), and (partially) Lazo (Fernandez et al., 2019). We were able to test Starmie on some of the data lakes (Binary, YADL Base, YADL 10k) though we did have to adapt and optimize certain sections of the discovery code that did not scale well to our environment. We were unable to run Starmie on the larger YADL 50k, and the less structured Open Data US. Starmie results are illustrated in Section 5. The modified Starmie repository is available here: https://github.com/rcap107/starmie, and we are contributing it upstream. These modifications are code optimizations and fixes for corner cases; they do not change the logic of Starmie.
Lazo relies on a client-server architecture that requires Elasticsearch v7. We were able to set up the backend environment and integrate the Lazo client in our pipeline, but we are unable to execute querying operations on our data lakes because the client database crashes due to out of memory errors. The wrappers we developed for testing Lazo are available in the main repository (https://github.com/rcap107/retrieve-merge-predict).
Additionally, for the retrieval step we considered ALITE (Khatiwada et al., 2022), KOIOS (Mundra et al., 2023), JOSIE (Zhu et al., 2019), and Saibot (Huang et al., 2023a). Each method requires substantial changes to the original codebase or very specific build configurations to run on datasets and environments different from those included in the original paper.
For the prediction pipeline, we consider ARDA (Chepurko et al., 2020) and Metam (Galhotra et al., 2023), focusing on Metam as it builds upon ARDA. Similarly to the retrieval case, Metam required substantial changes to be executed on data other than that in the paper and when tested in our environment it did not produce augmented tables. We implement some of the solutions suggested by Metam in our Stepwise Greedy Join selector, and we re-use the “Open Data US” data lake and “schools” base table in our experimental section. The modifications we made to the Metam code to interact with our pipeline are available at https://github.com/rcap107/metam.
We were not able to find an open repository for SilkMoth (Deng et al., 2017b), DeepJoin (Dong et al., 2023), Mileena (Huang et al., 2023b), or Kitana (Huang et al., 2023c).
Table 6: Recorded peak RAM usage for building and querying the different retrieval methods. All sizes are reported in MB. Runs that failed are marked as “NA”.
Case
Data Lake
Version
Exact
Matching
Minhash
Minhash
Hybrid
Starmie
Disk Binary 0.0 1.4 1.4 501.1
Open Data US 12.7 434 434 NA
YADL 10k 21.3 515 515 522.3
YADL 50k 28.2 2480 2480 NA
YADL base 21.3 462 462 522.4
RAM Build Binary 1196 226 226 19424
Open Data US 5466 3288 3286 NA
YADL 10k 688 3312 3290 87716
YADL 50k 1259 13697 13697 NA
YADL base 1061 2810 2810 38683
RAM Query Binary 255 216 216 4567
Open Data US 384 7789 10472 NA
YADL 10k 388 9138 12679 134869
YADL 50k 781 37332 37327 NA
YADL base 412 6784 8541 145437
C.2 Experimental setup
Table 7: Total equivalent compute hours, days, months and years required to run all the experiments. Single-CPU or single-GPU equivalent time: having a 32-CPU computer divides the time by 32.
Predictor Platform Total compute time
RidgeCV CPU 4y 3m 10d 7h
CatBoost CPU 1y 3m 29d 21h
ResNet GPU 5y 6m 23d 0h
RealMLP GPU 10y 7m 23d 3h
Total Both 21y 9m 26d 8h
We run our experimental campaign on a SLURM cluster, fixing the number of threads to 32. Nodes have at least 256GB of RAM. Experiments that involved NNs were run on nodes equipped with GPUs. Overall, preparing the retrieval methods required about 1050 CPU hours, while the experimental results required about 189k compute hours (about 21 years of equivalent CPU time); CatBoost and RidgeCV were run on CPU, ResNet and RealMLP on GPU. Table 7 reports the equivalent CPU (GPU) time across all experiments, broken by ML model; running experiments with 32 CPUs would reduce this runtime by about 32x.
Implementation details
The implementation of the pipeline is in Python [^8] Repository: https://github.com/rcap107/retrieve-merge-predict; Exact Matching, aggregation and join operations are implemented using Polars (Vink et al., 2024) as backend. For the Prediction step, RidgeCV is implemented with Scikit-learn (Pedregosa et al., 2011), CatBoost using the official library (Prokhorenkova et al., 2018), ResNet and RealMLP use pytabkit (Holzmüller et al., 2024).
We rely on the Python implementation of MinHash provided by the Datasketch package (Zhu et al., 2023). For Starmie, we optimize the original implementation so it can scale to work in our environment. We use the official DFS package. Data structures are stored by persisting on disk the ensembles for MinHash, the candidate ranking for each query column for Exact Matching, and the model checkpoints for Starmie; Hybrid MinHash relies on the MinHash data structures to work, so no additional storage is required for it.
Table 6 reports the size on disk of the data structures used for each method.
Retrieval We use a containment threshold of 0.2 for the preparation of the MinHash index, and clamp the number of candidates returned by each retrieval method to 30. These values were chosen to balance execution time and expected number of candidates given the distribution of containment encountered in the different data lakes (Figure 8). For Starmie, we use the default parameters defined in the original repository.
Selection We fix the number of Stepwise Greedy Join iterations to 30: this number is consistent with the number of candidates that are provided in the retrieval step. None of the other join selectors have parameters to tweak.
Prediction We fix the number of CatBoost iterations to 300; we stop training the model 10 iterations after the optimal metric has been detected; we set the L2 regularization coefficient to 0.01. As we are interested in evaluating the effect of each pipeline task rather than optimizing the prediction performance, we do not apply HPO to CatBoost.
We use the default parameters for RidgeCV as used in the scikit-learn implementation.
For RealMLP and ResNet we use the parameters that are set in the pytabkit package as they have been shown in Holzmüller et al. (2024) to be the “better defaults”.
C.3 Critical difference plot
We use critical difference plots to represent an overall ranking of the downstream prediction performance of all the configurations we considered. Specifically, the prediction metric is averaged over all base tables and data lakes, and the resulting value is used to rank each configuration.
Similarly to Figure 2 (and other Pareto plots), we present the same results after splitting them in retrieval method (Figure 11), join selector (Figure 11), and ML model (Figure 11).
Figure 9: Critical difference plot for the retrieval methods: Exact Matching, MinHash, Starmie, Hybrid MinHash.
Figure 10: Critical difference plot for the join selectors: Full Join, Stepwise Greedy Join, Highest Containment Join, Best Single Join.
Figure 11: Critical difference plot for the predictors: CatBoost, RidgeCV, RealMLP, ResNet.
C.4 Aggregated results
Table 8: Aggregated prediction results broken by base table and data lake. Starmie results are not reported here because it could not run on Open Data US and YADL 50k. As many ResNet and RidgeCV runs did not converge, we report a trimmed mean with cutoff 0.2.
Base Table Binary YADL Base YADL 10k YADL 50k Open Data US
Company Employees 0.086 $\pm$ 0.036 0.081 $\pm$ 0.043 0.071 $\pm$ 0.052 0.113 $\pm$ 0.048 0.011 $\pm$ 0.019
Housing Prices 0.16 $\pm$ 0.068 0.187 $\pm$ 0.077 0.143 $\pm$ 0.074 0.199 $\pm$ 0.092 0.221 $\pm$ 0.09
Schools NA NA NA NA 0.892 $\pm$ 0.137
2021 US Accidents 0.138 $\pm$ 0.065 0.178 $\pm$ 0.054 0.199 $\pm$ 0.076 0.150 $\pm$ 0.058 0.229 $\pm$ 0.028
US County Population 0.725 $\pm$ 0.203 0.078 $\pm$ 0.064 0.075 $\pm$ 0.074 0.112 $\pm$ 0.148 NA
2020 US Elections 0.374 $\pm$ 0.062 0.373 $\pm$ 0.115 0.358 $\pm$ 0.054 0.394 $\pm$ 0.045 0.400 $\pm$ 0.06
Figure 12: Pareto frontier analysis for the pipeline steps.
Figure 12 is prepared like Figure 2, however this version includes results from all data lakes and does not include Starmie as it could not run on Open Data and YADL 50k. Results are quite consistent with Figure 2: this means that the new data lakes (YADL 50k and Open Data US) do not alter the overall trends much.
In Figure 15 we report the fold vs fold difference in prediction score ($R^{2}$ and AUC) and relative execution time with respect to the reference configuration (Exact Matching, Best Single Join, aggregation Any and CatBoost) for retrieval (Figure 15 (a)), selection (Figure 15b), aggregation (Figure 15c), and prediction ( Figure 15d). Individual folds are reported as dots; color palettes depend on the data lake. This is an alternative way of representing the data in Table 3. Note that as we are comparing folds against folds, the difference is 0 when a particular parameter is the same as the reference. We choose Best Single Join as a reference for the selector because DFS could not run with Full Join and Stepwise Greedy Join.
Retrieval
Figure 15(a) shows that Exact Matching and Starmie have very similar performance (a median difference of 0.06%), with Starmie outperforming Exact Matching in some instances (this can also be observed in Figure 2 and Figure 11), which makes sense as the similarity metric used by Starmie combines Jaccard similarity with the cosine similarity of column embeddings. Base MinHash performs very poorly, with a median difference of -17.48% with respect to Exact Matching. This is not surprising and is likely caused by the lack of a candidate ranking combined with the presence of a candidate budget. Hybrid MinHash shows a marked improvement over base MinHash, gaining about 14% in median: this confirms that it is an effective strategy to address some of the shortcomings of the base method.
MinHash is faster than the others, despite the fact that the candidate budget $k$ is 30 for all methods. This is because, on average, the candidates retrieved by MinHash have a much lower containment than those proposed by the other methods (Figure 3a): due to the smaller amount of data to move and use for training the models, the runtime is shorter; furthermore, imperfect recall affects the performance of both MinHash and its hybrid variant.
While Hybrid MinHash appears to be faster than Exact Matching in the pipeline, it is important to note that re-ranking candidates incurs a non-negligible additional cost (Figure 15(a), Table 10).
Overall, the two methods based on precise ranking (Exact Matching and Starmie) outperform the methods based in approximate matching (MinHash and Hybrid MinHash), suggesting that the computational cost incurred in measuring exact containment does result in better prediction performance.
Selection
The choice of selector shows a clear effect when moving from single-table selectors (Highest Containment and Best Single Join) to multi-table selectors (Stepwise Greedy and Full Join), bringing a benefit of up to 2.2% in median (Figure 15(b)).
Stepwise Greedy Join is much slower compared to all other methods. This is not surprising, since this selector executes a join and trains a model in each iteration, then re-trains the model at the end of the fit step. In practice, while this added complexity brings an improvement of about 1.6% with respect to Best Single Join, this comes at the cost of a 2.9x increase in run time; even worse, this does not achieve the best performance as Full Join is both faster and slightly better in prediction performance. Highest Containment is significantly faster than all other solutions because it only re-ranks candidates before joining the top-$1$ candidate.
The difference in performance between the two single-table selectors (Best Single Join and Highest Containment) is likely due to the fact that, while Jaccard Containment is an indicator of a potentially good join, it is not sufficient for selecting the best candidate. For example, multiple candidates may have the same containment value, leading to ties. Given the lack of better information, Highest Containment breaks ties at random, thus it may select tables that are not relevant. Redundant data lakes like YADL 50k (Figure 8) are more affected by this problem. The significant difference between single- and multi-table selectors is explained by the learning model benefiting from an increased set of features: merging more than one table inherently results in a richer feature set.
Aggregation
Aggregation experiments do not include Full Join and Stepwise Greedy Join because DFS ran out of resources with those selectors. In fact, while DFS brings some benefit in its generation of new features (up to 2.46% in median wrt Any), it is also extremely slow (up to 6x wrt the simpler method), and this was in a simpler case where only one table was joined at a time. With multi-table selectors, this problem became even more noticeable and prevented us from testing the multi-table selectors. This very large memory cost is consistent with Cvetkov-Iliev et al. (2023).
For what concerns the simpler methods, Mean is slightly better than Any, and has a slightly longer runtime; we did not observe major differences between the two methods in practice.
Predictor
The performance difference between CatBoost and all other methods already highlighted in Figure 2 and Figure 11 is confirmed in Figure 15(d), which shows CatBoost outperforming the second best model in RealMLP by about 21.2%, and the worse NN model in RidgeCV by 26%. Interestingly, CatBoost is also faster than all other methods by a significant amount, likely owing to the optimized preprocessing of categorical variables. RealMLP is the better performing NN model, and has shown a good resilience in the very challenging environment that we are considering; however, it is much slower than all other methods, even when run on GPUs, requiring up to 32x as much time as CatBoost.
Different Lakes
Table 8 reports the trimmed average of the prediction performance across all configurations, broken by base table and data lake. We observe experimentally that many RidgeCV and ResNet runs do not converge, leading to very long run times and extremely large negative values for $R^{2}$. For this reason, we use a trimmed mean with a cutoff of 0.2. Indeed, the results are so poor because the mean across predictors is lowered substantially by the performance of the parametric models. Internal tables could were only run on the data lake they were sampled from, and other cells are reported as “NA”.
Despite its lower containment compared to YADL-based data lakes (Figure 3a and Figure 8), Open Data US shows good results, outperforming the YADL variants in some cases: this may be due to the smaller fraction of nulls, and larger tables on average compared to YADL.
Prediction performance was particularly good on US demographic-adjacent tables, and on the internal dataset “Schools”, which achieved perfect classification performance in most cases – also visible in Figure 3b, where highest containment leads to best performance.
YADL’s internal table (US County Population) exhibits a similar, though not quite as extreme, behavior to that of Schools. This can be explained by the fact that internal tables, i.e., tables that are sampled from the data lake itself, tend to have copies in the data lake, which are likely to contain information that is correlated with the prediction task. This highlights the specific behavior of internal tables, which are routinely used in the experimental campaigns in the literature.
The different prediction performance obtained with each YADL variant suggest that the structure of each data lake impacts the actual prediction performance.
C.5 Smart aggregation brings some benefit at a major cost
Figure 13: Aggregation trade-offs relative to run time.
Figure 14: Aggregation trade-offs relative to peak RAM.
While containment impacts heavily both prediction performance and execution time, aggregation has a similar impact on the execution time, without the same degree of improvement for the prediction performance (Figure 15c).
Aggregating values always leads to a loss of information: in exchange for a larger cost, complex aggregation methods that preserve more information (DFS) or that replace values with better representatives of a group (mean) should lead to better prediction performance. To an extent, this is what we observe: DFS does improve the downstream performance by a decent amount, however this is at a major computational cost. This problem is exacerbated by the fact that aggregation must be performed whenever a join is executed at any point in the pipeline, including joins executed during join selection. The result is a compounding slow-down of the entire pipeline.
The similar performance of “any” and “mean” can be explained by the execution of joins on key-like columns so that values in these columns tend to be mostly distinct, thus producing the same aggregated value for both “mean” and “any”. The slight difference in performance and runtime between the two methods may be attributed to the large fraction of missing values in some of the data lakes: the “mean” method is biased towards selecting the most frequent value, which may be a null. When that happens, more values than needed become nulls, reducing the amount of features and introducing a slight speedup in the overall training.
The relatively unsatisfying performance of DFS is also explained by the fact that we are not fully exploiting its capabilities. We consider only join depth-1 chains: at each aggregation step, we join the base table with an additional table, rather than leveraging the recursive generation of features provided by DFS. As a result, DFS is not as effective at generating features as it would be with deeper join paths (Cvetkov-Iliev et al., 2023) and may benefit from integrating join path discovery systems (Deng et al., 2017a).
Figure Figure 13 and Figure 14 confirm that, while DFS improves the prediction performance, its overall cost remains problematic.
Figure 15: Experimental results across all data lakes, comparing the performance difference between evaluated methods in different steps of the pipeline. The median difference is reported on the right of each plot. (a) compares join selectors, (b) aggregation methods, and (c) ML models. In all cases, results are relative to the “average method”. Starmie results are not reported in this plot as we could not run it on the larger data lakes.
C.6 Distribution of the containment for each retrieval method
To understand the success of the different retrieval methods, we look at the containment of the joins that they suggest. Figure 16 reports top-200 containment. It highlights that, even when retrieving a large number (200) of candidates, the average containment of MinHash is very low compared to the other methods: for larger datasets, MinHash returns thousands of candidates of which only a fraction are selected. The lack of an internal ordering means that high-containment candidates are likely to be missed (Section 4). While MinHash has a threshold that can be tweaked to reduce the number of candidates that are retrieved, we observe that recall drops sharply at high thresholds. Hybrid Minhash has some success in mitigating the problem, as it improves the average containment and the overall downstream performance (Figure 2 left).  [^9] The performance of MinHash on Binary is an artefact: the method could only retrieve about 6 candidates on average compared to the 30 found by the other two methods, thus values are averaged across fewer candidates with higher containment. We were unable to obtain the containment results for Starmie on YADL 50k and Open Data US.
Figure 16: Distribution of containment in the top-200 candidates obtained by each retrieval method on different data lakes.
C.7 Execution time breakdown
Figure 17: Breakdown of where time is spent for each selector. On the right we report the average run time of each selector.
Figure 18: Relative prediction performance (median on the right) obtained selecting the (top-$1$) candidate according to the ranking provided by the retrieval methods. YADL 50k and Open Data are not included.
Figure 18 breaks down where the time is spent by each join selector, with the corresponding average total time on the right. Prepare tracks the time spent building data structures (including any ranking operation), and loading data. Train(model) and Predict(model) track the time spent inside the ML model for training and prediction, respectively. Finally, Train(join) and Predict(join) track the time spent executing a merge operation, combining join and aggregation. Results are aggregated over all experiments with “first” as aggregation. The time distribution follows our expectations: most of the time is spent fitting models, with time spent joining (and aggregating) in second place. Highest Containment Join spends a relatively long time in the “Prepare” step due to the need to re-rank candidates before joining: since the join and train steps involve only one table, they are faster in comparison. Stepwise Greedy Join has a similar re-ranking step, however training the models in each iteration dominates the other steps. Full Join merges all candidates at the same time, then trains a single model on the result: this explains how the fraction of time spent joining is larger than in other methods. Top-$1$ Full Join is a reference of how long it would take to simply join one candidate and train on that, without any additional operation: as expected, it is very fast.
C.8 Effectiveness of top-$1$ join with different retrieval methods
Figure 18 reports the prediction performance across different data lakes obtained by joining each base table on the top-$1$ candidate according to each retrieval method’s internal ranking (here, Exact is equivalent to top-$1$ Highest Containment Join). Unsurprisingly, MinHash is far worse than all other methods (up to 13.5% in median) due to the lack of an inherent ordering, and budget constraints; Hybrid MinHash successfully mitigates some of the issues of MinHash, however the imperfect retrieval performance penalizes it. It appears that the first candidate retrieved by Starmie is not as good as what is selected by Exact Matching. We observe that the performance of Starmie drops on YADL 10k: this is likely due to the larger width of the candidate tables, which affects the cosine similarity, and thus worsens the ranking. However, the situation improves for multi-table selectors, suggesting that this issue is mitigated by selecting multiple candidates at once.
Table 9: Statistics for the base tables.